CO 3) A random selection from a deck of cards selects one card. What is the probability of selecting an ace?


0.077


0.50

0.250

0.025
Answer #1

There are 4 aces in a deck of 52 cards.

So, the probability of selecting a ace from a deck of cards = 4 / 52 = 1/13 = 0.07692308

\approx0.077

The correct answer is 0.077